OpenStudy (anonymous):

I don't want the answer, just the steps to get there. Thanks for your help in advance. A wrench 30 cm long lies along the positive y-axis and grips a bolt at the origin. A force is applied in the direction <0, 3, -4> at the end of the wrench. Find the magnitude of the force needed to supply 100N/m of torque to the bolt.

hero (hero):

Post this in Physics

OpenStudy (anonymous):

Its Math 231 (Motion in Space: Velocity and Acceleration)

OpenStudy (amistre64):

not saying it aint math, but a post up in physics would be useful as well since its a physics application :)

OpenStudy (anonymous):

Alright, I'll give it a try.

OpenStudy (amistre64):



OpenStudy (anonymous):

@amistre64 "force is applied in the direction <0, 3, -4> at the end of the wrench." [considering only y-z plane] this would mean that vector F passes through (3, -4) and (30, 0) right?

OpenStudy (amistre64):

it would appear to me that the magnitude of the downward force, times the length of the wrench Vaide, i believe so, yes

OpenStudy (amistre64):

well, the vector need not pass thru 3,-4; but its direction from the end is such

OpenStudy (anonymous):

Initial point (0i, 30j, 0k) < 0, 3, -4> = Velocity? Acceleration? If Acceleration: F=ma -> F/m = a -> 100N/m = < 0, 3, -4>? Then anti-derivative of < 0, 3, -4> to Velocity -> then take the magnitude of the Velocity?

OpenStudy (anonymous):

so we can find the slope of F which will be equal to tan(theta). from this we can know theta. so, T = F.s. sin(theta)

OpenStudy (amistre64):

id have to read up on torque to be certain :)

OpenStudy (anonymous):

@MCouch don't proceed that way. use the formula for torque.

OpenStudy (amistre64):

torque = distance cross force

OpenStudy (anonymous):

here's a brush-up: torque = r x F

OpenStudy (anonymous):

so, T = F.r.sin(theta) we know T, r, theta. find F.

OpenStudy (amistre64):

x 0 0 x = -120 y 30 3 y = 0 z 0 -4 z = 0 as is; you seem to have 120 somethings

OpenStudy (amistre64):

to scale the force use 0s 3s and -4s is my idea x 0 0s x = -120s y 30 3s y = 0 z 0 -4s z = 0 where 120s = 100; or s= 100/120, 5/6

OpenStudy (anonymous):

@amistre64 That makes more sense to me @Vaidehi09 We have not gone over torque in that way yet...

OpenStudy (amistre64):

the force direction creates a 3,4,5 right triangle; so the actual force might be 5s

OpenStudy (anonymous):

oh? well, then use whichever way suits u!

OpenStudy (amistre64):

i used this as a stepping stone http://www.physics.uoguelph.ca/tutorials/torque/Q.torque.intro.html

OpenStudy (amistre64):

still not quite sure how to convert that to Nm

OpenStudy (amistre64):

id imagine 30cm = .3 meters which would have to be used in the stuff <0,.3,0> instead

OpenStudy (anonymous):

In my notes from a previous problem we did in class, we had 20N = 4kg x a we then converted it into a(t) = < 0, 0, 5> b/c it was stated the force was only in the positive z-axis We then did the anti-derivative of a(t) to find v(t)

OpenStudy (amistre64):

im not sure how to convert those to at and vt :/

OpenStudy (anonymous):

Well thanks for the help thus far

OpenStudy (amistre64):

one newton/meter is equal to the torque applied by one newton on a lever of one meter is what the wiki says ...

OpenStudy (amistre64):

if we shorten the lever, we lessen the torque right? in any event, im going with distance in meters crossed with the force vector :)

OpenStudy (amistre64):

is there an answer key we can chk? cause i think i come up with about 416

OpenStudy (anonymous):

No answer key :( and i'm trying to get in touch with a class mate. And this is" MATH 231 Calculus of Several Variables (2) Analytic geometry in space; partial differentiation and applications. " If that helps

OpenStudy (amistre64):

my thoughts are: i define torque as the amount of pressure at the end of the lever; that pressure is calculated as length * perpendicular force. a NewtonMeter is a length of 1 meter, with a force of ! newton 30 cm = .3m ; the perp force is the "4" portion of the force vector; and the actual force of the vector is the hypotenuse of the 3,4,5 triangle it creates (the 5 part) 100 = .3 * 4s 100 = 1.2s s = 1000/12 5s = 5000/12 = 416 2/3 not that its right or anything mind you :)

OpenStudy (anonymous):

I will keep that in mind maybe it is right :) But either way I'm going to take a break and hopefully my classmate will respond. I will also continue to dig through my textbook and notes for something that will give me heads or tails of this...

OpenStudy (anonymous):

Thanks again.

OpenStudy (amistre64):

good luck ;)

OpenStudy (anonymous):

@amistre64 You were correct with 416.5. You loosely follow F=ma or F/m =a F/m = 100N/m and a = Starting point < 0, -.3, 0> cross directional vector < 0, 3, -4> a = 83.3 and you have to do <0a, 3a, -4a> to get the vector and then take the magnitude of it.